Multiply by fractions (fraction of)

Multiplying by a fraction always means finding that fraction of the quantity

Here it is half of 1/2 , or 1/2× 1/2

1. What is a half of 1/2?( You can see it by putting the quarter strip next to the half strip)



1
Expert's answer
2021-06-29T14:43:23-0400

"\frac{1}{2}\times\frac{1}{2}\\"


"=\frac{1\times1}{2\times 2}"


"=\frac{1}{4}"